On Monday, 12 people working together finished a task in 3 hours. On Tuesday, the task had to be repeated, but some of the 12 people were gone. If the task was completed in 4 hours, how many people were gone?
Answer by ankor@dixie-net.com(22740)

On Monday, 12 people working together finished a task in 3 hours.
Find the no. of man-hrs required to do the job
12 *3 = 36 man-hrs
On Tuesday, the task had to be repeated, but some of the 12 people were gone. If the task was completed in 4 hours, how many people were gone?
let m = no. of men gone
4(12-m) = 36
48 - 4m = 36
-4m = 36 - 48
-4m = -12
m = -12/-4
m = +3 men were gone